# Relative equilibria for Scaling Symmetries and Central Configurations

### v1

@@ Abstract
In this paper, we explore scaling symmetries within the framework of symplectic geometry. We focus on the action $\Phi$ of the multiplicative group $G=\mathbb{R}^{+}$ on exact symplectic manifolds $(M,\omega,\theta)$, with $\omega=-d\theta$, where $\theta$ is a given primitive one-form. Extending established results in symplectic geometry and Hamiltonian dynamics, we introduce conformally symplectic maps, conformally Hamiltonian systems, conformally symplectic group actions, and the notion of conformal invariance. This framework allows us to generalize the momentum map to the conformal momentum map, which is crucial for understanding scaling symmetries. Additionally, we provide a generalized Hamiltonian Noether’s theorem for these symmetries.
@@ 1.1 Background and Motivation
Furthermore, by leveraging the conserved quantities encoded in the momentum map through the process of symplectic reduction [19], one can reduce the dimensionality of the phase space, leading to a new, lower-dimensional symplectic manifold known as the reduced space. The dynamics on this reduced space are governed by the reduced Hamiltonian equations. Relative equilibria can then be viewed as equilibrium points of the reduced Hamiltonian equations on the reduced space.
@@
Recently, however, Bravetti, Jackman, and Sloan [5] have shown that a symplectic Hamiltonian system with a scaling symmetry can be reduced to a contact Hamiltonian system with one fewer degree of freedom. This reduction utilizes a so-called scaling function, which plays a role somewhat analogous to that of equivariant momentum maps in symplectic reduction. An interesting feature of this approach is that it allows for various choices of scaling functions to obtain the reduced equations. However, in practice, there may be no obvious choice of a scaling function. In this paper, we address this issue by identifying a distinguished scaling function, which we call the conformal momentum map, that generalizes the momentum map of symplectic reduction.
@@
The approach to defining central configurations applies to the Kepler and $n$-body problems on surfaces of revolutions that admit scalings, the Anisotropic Kepler problem and to the Kepler and two-body problem on the Heisenberg group. In the latter case the application is somewhat trivial as the action of the scaling symmetry is symplectic and the Hamiltonian is conformally invariant in that context [26]. Unfortunately, however, the approach adopted in this study associates the definition of central configurations with scaling symmetries, making it infeasible to extend this definition to spaces of constant curvature. In fact, although the sphere and the hyperbolic plane are homogeneous and isotropic, they do not permit dilations, thereby preventing the application of our central configuration definition to these geometries.
@@ 1.2 Summary of Main Results and Outline of the Paper
In this paper, we study systems with scaling symmetries within the framework of symplectic geometry. To effectively capture these symmetries, we consider the action $\Phi$ of the multiplicative group $G=\mathbb{R}^{+}$ on an exact symplectic manifold $(M,\omega)$, where $\omega=-d\theta$.
@@
We specialize our formulas for the momentum map to the case of the cotangent bundle. Given an action $\Psi$ of $G=\mathbb{R}^{+}$ on $Q$, we define a scaled cotangent lifted action on $T^{*}Q$. In this context, the conformal momentum map is given by:
@@ 2 Conformally Hamiltonian systems
$d(H\circ f)(z)\cdot v=dH(f(z))\cdot[Tf\cdot v]=dH(f(z))\cdot[f_{\ast}v]=\omega(f(z))(X_{H}(f(z)),f_{\ast}v)$
@@ 3 Cotangent Bundles and Scaled Cotangent Lift
In many examples or applications, the phase space is the cotangent bundle $M=T^{\ast}Q$ of a configuration space $Q$. The cotangent bundle $M=T^{\ast}Q$, equipped with the projection map $\pi_{Q}:M\to Q$, can be endowed with the one-form $\theta_{0}$ defined by
$\langle\theta_{0}(\beta),v\rangle=\langle\beta,T\pi_{Q}\cdot v\rangle=\left\langle(T\pi_{Q})^{\ast}\beta,v\right\rangle,$
[…]
@@
While the primary goal of this paper is not to describe dynamical systems with friction, it is noteworthy that the formalism of conformal Hamiltonian systems can be used to describe such systems, as it is illustrated in the following example.
@@
which yields the same Newton’s equation as the Hamiltonian $H$. Moreover, note that in this case $X_{F}^{c}=(p,-bp-\frac{\partial U}{\partial q})$, $\theta_{0}=p\,dq$, and $a=-b$. Hence,
@@
We will now introduce an important method for generating conformal symplectic transformations on cotangent bundles. To achieve this, we begin by providing the following definition, which extends the concept of cotangent lifts.
@@
Let $Q_{1}$ and $Q_{2}$ be manifolds and $f:Q_{1}\to Q_{2}$ a diffeomorphism. Then the scaled cotangent lift $\hat{f}^{\lambda}:T^{\ast}Q_{1}\to T^{\ast}Q_{2}$ of $f$ with parameter $\lambda\in\mathbb{R}$ to the cotangent bundle is defined by
$\langle\hat{f}^{\lambda}(\alpha),v\rangle$ $=\left\langle\alpha,\lambda(Tf^{-1}(v))\right\rangle=\lambda\left\langle\alpha,Tf^{-1}(v)\right\rangle$
where $q\in Q_{1}$, $\alpha\in T_{q}^{\ast}Q_{1}$, $v\in T_{f(q)}Q_{2}$, $Tf^{-1}$ is the tangent map (or tangent lift) of $f^{-1}$.Clearly, $\hat{f}^{\lambda}=\lambda\hat{f}$, where $\hat{f}$ is the standard cotangent lift.
@@
Notice that $\hat{f}^{\lambda}=\lambda\hat{f}$, where $\hat{f}$ is the cotangent lift of the diffeomorphism $f$. It is well known that $\hat{f}^{\ast}\theta_{Q_{2}}=\theta_{Q_{1}}$ (see Theorem 5.11 in [18]). Let $\Lambda$ be the function defined by $\Lambda(z)=\lambda z$, so that we can write $\hat{f}^{\lambda}=\Lambda\circ\hat{f}=\lambda\hat{f}$. Then, we have
$(\hat{f}^{\lambda})^{\ast}\theta_{Q_{2}})(v)$ $=((\Lambda\circ\hat{f})^{\ast}\theta_{Q_{2}})(v)$
$=(\hat{f}^{\ast}(\Lambda^{\ast}\theta_{Q_{2}}))(v)\quad\mbox{by the properties of pullback}$
$=(\hat{f}^{\ast}\theta_{Q_{2}})(\Lambda_{\ast}v)\quad\mbox{by thedefinition of pullback}$
$=(\hat{f}^{\ast}\theta_{Q_{2}})(\lambda v)\quad\mbox{by the definition of $\Lambda$ and pushforward}$
$=\lambda(\hat{f}^{\ast}\theta_{Q_{2}})(v)\quad\mbox{by linearity}$
$=\lambda\theta_{Q_{1}}(v)\quad\mbox{since $\hat{f}^{\ast}\theta_{Q_{2}}=\theta_{Q_{1}}$},$
for every $v\in T^{\ast}Q_{1}$. This completes the proof. ∎
@@
Since the exterior derivative commute with the pull-back, we get
@@
In coordinates $(\mathbf{q},\mathbf{p})$, the scaled cotangent lift can be explicitly written as:
$\hat{f}^{\lambda}(\mathbf{q},\mathbf{p})=\left(f(\mathbf{q}),\lambda(Df(\mathbf{q}))^{-T}\cdot\mathbf{p}\right).$ (2)
Here, $Df(\mathbf{q})$ denotes the Jacobian matrix of $f$ evaluated at $\mathbf{q}$, and $(Df(\mathbf{q}))^{-T}$ represents the inverse transpose of this Jacobian matrix. The explicit expression in coordinates simplifies the application of the scaled cotangent lift in concrete problems.
@@ 4 Conformally Symplectic Lie Group Actions
is a flow on $M$, where $\exp:\mathfrak{g}\to G$ is the exponential map. [^1] We will use $\exp(x)$ to denote the exponential map and $e^{x}$ to denote the usual exponential function of real numbers. To each element $\xi\in\mathfrak{g}$, there is an associated vector field on $M$, which we denote by $\xi_{M}$, and call the infinitesimal generator of the action corresponding to $\xi$. This vector field is defined as follows:
@@
We now introduce some important concepts for this work, namely, conformally symplectic actions and scaling symmetries. The next definition generalizes the idea of symplectic action.
@@
Let $(M,\omega,\theta)$ an exact symplectic manifold. An action $\Phi:\mathbb{R}^{+}\times M\to M$ of the multiplicative group $G=\mathbb{R}^{+}$ on $M$ is said to be conformally symplectic with parameter $c$ if there exist $c\in\mathbb{R}$ such that
$\Phi_{g}^{\ast}\omega=g^{c}\omega,$
for all $g\in G=\mathbb{R}^{+}$ In the particular case in which $c=0$ the action is symplectic.
Since this paper primarily focuses on the multiplicative group $G=\mathbb{R}^{+}$, it is natural to specialize this definition to this group. The Lie algebra of $G=\mathbb{R}^{+}$ can be identified with $\mathbb{R}$, in such a way that the exponential map is just the usual exponential of real numbers. So here $\mathfrak{g}=\mathbb{R}$. Then for $\xi\in\mathfrak{g}=\mathbb{R}$ the map $\Phi_{e^{\tau\xi}}:M\to M$ is a flow on $M$. The infinitesimal generator of the action corresponding to $\xi$ is
@@
Then if $\Phi_{g}^{\ast}\omega=g^{c}\omega$ we have
$\mathcal{L}_{\xi_{M}}\omega=\left.\frac{d}{d\tau}\right|_{\tau=0}\phi_{\tau}^{\ast}\omega=\left.\frac{d}{d\tau}\right|_{\tau=0}\Phi_{e^{\tau\xi}}^{\ast}\omega=\left.\frac{d}{d\tau}\right|_{\tau=0}e^{(\tau c\xi)}\omega=(c\xi)\,\omega$
We call a group action for which $\mathcal{L}_{\xi_{M}}\omega=(c\xi)\,\omega$ infinitesimally conformally symplectic. The next definition generalizes the notion of invariant function.
@@
Let $(M,\omega,\theta)$ an exact symplectic manifold and $F:M\to\mathbb{R}$ a function on $M$. Suppose the action $\Phi:\mathbb{R}^{+}\times M\to M$ of the multiplicative group $G=\mathbb{R}^{+}$ on $M$ is conformally symplectic with parameter $c$. Then, we say that $F$ is conformally invariant if there exist $b\in\mathbb{R}$ such that
$\Phi_{g}^{\ast}F=g^{b}F$
for all $g\in G=\mathbb{R}^{+}$. If $b=0$ we say that the function is invariant.
If $G=\mathbb{R}^{+}$ and $\xi_{M}$ is the infinitesimal generator corresponding to $\xi\in\mathfrak{g}=\mathbb{R}$. Then
$\mathcal{L}_{\xi_{M}}F=\left.\frac{d}{d\tau}\right|_{\tau=0}\Phi_{e^{\xi\tau}}^{\ast}F=\left.\frac{d}{d\tau}\right|_{\tau=0}e^{b\xi\tau}F=(b\xi)F.$
We refer to this property as infinitesimal conformal invariance. This concept extends the well-known notion of infinitesimal invariance (see [17] for the classical definition). According to the terminology used in [5], a function $F$ that exhibits this property is termed a scaling function.
@@
Let $(M,\omega,\theta,H)$ a Hamiltonian system on an exact symplectic manifold, and let $\Phi:\mathbb{R}^{+}\times M\to M$ be an action of the multiplicative group $G=\mathbb{R}^{+}$ on $M$. We say that $\Phi$ is a scaling symmetry if $\Phi$ is conformally symplectic and $H$ is conformally invariant. In other words $\Phi$ is a scaling symmetry if there exist $c,b\in\mathbb{R}$ such that
$\Phi_{g}^{\ast}\omega=g^{c}\omega,\quad\Phi_{g}^{\ast}H=g^{b}H$
for all $g\in\mathbb{R}^{+}$.
If $G=\mathbb{R}^{+}$ and if $\xi_{M}$ is an infinitesimal generator of the scaling symmetry corresponding to $\xi\in\mathfrak{g}=\mathbb{R}$ then
@@
The following proposition, which will be useful in our discussion of relative equilibria in Section 6, is analogous to Corollary 2.9 but pertains to scaling symmetries rather than diffeomorphisms. Its proof follows the same approach as that of Proposition 2.8 (see also [23]).
@@
Let $(M,\omega,\theta,H)$ a Hamiltonian system on an exact symplectic manifold, and let $\Phi:\mathbb{R}^{+}\times M\to M$ a scaling symmetry. Then
$\Phi^{\ast}_{g}X_{H}=g^{b-c}X_{H}.$

### v2

@@ Abstract
In this paper, we explore scaling symmetries within the framework of symplectic geometry. We focus on the action $\Phi$ of the multiplicative group $\mathbb{R}^{+}$ on exact symplectic manifolds $(M,\omega,\theta)$, with $\omega=-d\theta$, where $\theta$ is a given primitive one-form. Extending established results in symplectic geometry and Hamiltonian dynamics, we introduce conformally symplectic maps, conformally Hamiltonian systems, conformally symplectic group actions, and the notion of conformal invariance. This framework allows us to generalize the momentum map to the conformal momentum map, which is crucial for understanding scaling symmetries. Additionally, we provide a generalized Hamiltonian Noether’s theorem for these symmetries.
@@ 1.1 Background and Motivation
Furthermore, by utilizing the conserved quantities encoded in the momentum map through the process of symplectic reduction [18], one can reduce the dimensionality of the phase space, leading to a new, lower-dimensional symplectic manifold known as the reduced space. The dynamics on this reduced space are governed by the reduced Hamiltonian equations, and relative equilibria can then be viewed as equilibrium points of these reduced equations.
@@
Bravetti, Jackman, and Sloan [3] have recently shown that a symplectic Hamiltonian system with scaling symmetry can be reduced to a contact Hamiltonian system with one fewer degree of freedom. This reduction depends on a scaling function, which plays a role similar to equivariant momentum maps in symplectic reduction. A key feature of their approach is its flexibility, allowing for various choices of scaling functions to obtain the reduced equations. In systems on cotangent bundles with a Riemannian metric on the configuration space, there is often a natural choice for a scaling function, something that is implicitly apparent in [3]. For instance, in the two-body problem, a power of the configuration variable is a natural choice, while in the $n$-body system, a power of the moment of inertia serves this purpose. In this paper, we introduce a new distinguished scaling function, which we refer to as the conformal momentum map, extending the notion of the momentum map in symplectic reduction.
@@
The approach for defining central configurations introduced in this paper applies to various contexts, including the Kepler and $n$-body problems on surfaces of revolution that admit scaling symmetries, the Anisotropic Kepler problem, and the Kepler and two-body problem on the Heisenberg group. However, in the case of the Heisenberg group, the general theory is not fully utilized, as the scaling symmetry is symplectic, even though the Hamiltonian is conformally invariant [25]. Furthermore, note that the results from Section 7 cannot be used for the Kepler and two-body problem on the Heisenberg group, as the configuration space in these cases is endowed with a sub-Riemannian metric rather than a Riemannian one.
Unfortunately, however, the approach adopted in this study associates the definition of central configurations with scaling symmetries, making it infeasible to extend this definition to spaces of constant curvature. In fact, although the sphere and the hyperbolic plane are homogeneous and isotropic, they do not permit dilations, thereby preventing the application of our central configuration definition to these geometries.
@@ 1.2 Summary of Main Results and Outline of the Paper
In this paper, we study systems with scaling symmetries within the framework of symplectic geometry. To effectively capture these symmetries, we consider the action $\Phi$ of the multiplicative group $\mathbb{R}^{+}$ on an exact symplectic manifold $(M,\omega)$, where $\omega=-d\theta$.
@@
We specialize our formulas for the momentum map to the case of the cotangent bundle. Given an action $\Psi$ of $\mathbb{R}^{+}$ on $Q$, we define a scaled cotangent lifted action on $T^{*}Q$. In this context, the conformal momentum map is given by:
@@ 2 Conformally Hamiltonian systems
Remark 2.6.
Note that the concept of a locally conformally Hamiltonian vector field seems to be closely related to the notion of dynamical similarity, introduced by David Sloan in [28].
@@
$d(H\circ f)(z)\cdot v=dH(f(z))\cdot[df(z)\cdot v]=dH(f(z))\cdot[f_{\ast}(z)v]=\omega(f(z))(X_{H}(f(z)),f_{\ast}v)$
@@ 3 Cotangent Bundles and Scaled Cotangent Lift
In many applications, especially in classical mechanics, the phase space is often modeled as the cotangent bundle $M=T^{\ast}Q$ of a configuration space $Q$. The cotangent bundle $M$ is naturally equipped with a canonical 1-form, denoted $\theta_{0}$.
At any point $z=(q,p)\in M$, where $q\in Q$ and $p\in T_{q}^{*}Q$, the natural projection map $\pi:M\to Q$ takes a point $z$ in the cotangent bundle to its corresponding point $q$ in the configuration space. The differential of this projection at $z$, denoted $d\pi_{z}:T_{z}M\to T_{q}Q$, maps a tangent vector $v\in T_{z}M$ to a tangent vector $d\pi_{z}(v)\in T_{q}Q$.
The 1-form $\theta_{0}$, evaluated at $z$ and acting on $v\in T_{z}M$, is defined as:
$(\theta_{0})_{z}(v)=\langle p,d\pi_{z}(v)\rangle,$
where $p\in T_{q}^{*}Q$ is the covector at the base point $q$, and $\langle p,d\pi_{z}(v)\rangle$ represents the natural pairing between the covector $p$ and the tangent vector $d\pi_{z}(v)$.
Alternatively, $(\theta_{0})_{z}$ can be expressed as:
$(\theta_{0})_{z}=(d\pi_{z})^{*}p,$
where $(d\pi_{z})^{*}:T_{q}^{*}Q\to T_{z}^{*}M$ is the transpose of the map $d\pi_{z}:T_{z}M\to T_{q}Q$.
The one-form $\theta_{0}$ is called the Liouville one-form or canonical one-form. The exterior derivative $\omega_{0}=-d\theta_{0}$ is a symplectic form on $M$. The forms $\theta_{0}$ and $\omega_{0}$ are known as the canonical forms on $M=T^{\ast}Q$. In this paper, we will always consider $T^{\ast}Q$ with the canonical forms, treating $T^{\ast}Q$ as the exact symplectic manifold $(T^{\ast}Q,\omega_{0},\theta_{0})$.
@@
Although the primary focus of this paper is not the study of dynamical systems with friction, it is worth noting that the formalism of conformal Hamiltonian systems can be effectively applied to describe such systems, as demonstrated in [21] (see also [29]) and illustrated in the following example.
@@
which yields the same Newton’s equation as the Hamiltonian $H$. Moreover, since in this case $X_{F}^{c}=(p,-bp-\frac{\partial U}{\partial q})$, $\theta_{0}=p\,dq$, and $c=-b$, then by Proposition 2.3 we have:
@@
We now introduce an important method for generating conformal symplectic transformations on cotangent bundles by generalizing the concept of cotangent lifts.
Let $Q_{1}$ and $Q_{2}$ be manifolds, and let $f:Q_{1}\to Q_{2}$ be a diffeomorphism. Recall that the cotangent lift $\hat{f}:T^{*}Q_{1}\to T^{*}Q_{2}$ is defined as follows:
For each point $z_{1}=(q_{1},p_{1})\in T^{*}Q_{1}$, the cotangent lift $\hat{f}$ maps $z_{1}$ to a point $z_{2}=(q_{2},p_{2})\in T^{*}Q_{2}$, where:
$z_{2}=\hat{f}(z_{1})=\left(f(q_{1}),(df^{-1}_{f(q_{1})})^{*}(p_{1})\right),$
where
  • $df^{-1}_{f(q_{1})}:T_{f(q_{1})}Q_{2}\to T_{q_{1}}Q_{1}$ is the differential of the inverse map $f^{-1}$ evaluated at $f(q_{1})$,
  • $(df^{-1}_{f(q_{1})})^{*}:T_{q_{1}}^{*}Q_{1}\to T_{f(q_{1})}^{*}Q_{2}$ is the transpose of $df^{-1}_{f(q_{1})}$, which maps the covector $p_{1}\in T_{q_{1}}^{*}Q_{1}$ to a covector in $T_{f(q_{1})}^{*}Q_{2}$.
To extend this idea, we introduce a scaled version of the cotangent lift that includes a scaling factor:
@@
Let $Q_{1}$ and $Q_{2}$ be manifolds, and let $f:Q_{1}\to Q_{2}$ be a diffeomorphism. The scaled cotangent lift $\hat{f}^{\lambda}:T^{*}Q_{1}\to T^{*}Q_{2}$, with a scaling parameter $\lambda\in\mathbb{R}$, is defined by:
$\hat{f}^{\lambda}(z_{1})=\left(f(q_{1}),\lambda\cdot(df^{-1}_{f(q_{1})})^{*}(p_{1})\right),$
for each $z_{1}=(q_{1},p_{1})\in T^{*}Q_{1}$.
In coordinates $(\mathbf{q},\mathbf{p})$ on $T^{\ast}Q_{1}$, the scaled cotangent lift is expressed as:
$\hat{f}^{\lambda}(\mathbf{q},\mathbf{p})=\left(f(\mathbf{q}),\lambda(Df(\mathbf{q}))^{-T}\cdot\mathbf{p}\right),$ (2)
where $Df(\mathbf{q})$ is the Jacobian matrix of $f$ evaluated at $\mathbf{q}$, and $Df(\mathbf{q})^{-T}$ denotes its inverse transpose. This explicit coordinate expression simplifies the application of the scaled cotangent lift in concrete problems by providing a direct representation of the transformation.
@@
We aim to show that $((\hat{f}^{\lambda})^{*}\theta_{Q_{2}})_{z_{1}}=\lambda(\theta_{Q_{1}})_{z_{1}}$ for any point $z_{1}=(q_{1},p_{1})\in T^{*}Q_{1}$.
We will use the following facts in the proof:
  • Let $z_{2}=\hat{f}^{\lambda}(z_{1})=(q_{2},p_{2})\in T^{*}Q_{2}$. By the definition of the scaled cotangent lift $\hat{f}^{\lambda}$, we have:
    $q_{2}=f(q_{1}),\quad\text{and}\quad(df_{q_{1}})^{\ast}p_{2}=\lambda p_{1}.$
  • The canonical 1-forms $\theta_{Q_{1}}$ and $\theta_{Q_{2}}$ are defined as:
    $(\theta_{Q_{1}})_{z_{1}}=(d\pi_{1})^{\ast}_{z_{1}}p_{1},\quad(\theta_{Q_{2}})_{z_{2}}=(d\pi_{2})^{\ast}_{z_{2}}p_{2},$
    where $\pi_{i}:T^{*}Q_{i}\to Q_{i}$ ( $i=1,2$) are the natural projections $\pi_{i}(q_{i},p_{i})=q_{i}$.
  • For any $z_{1}=(q_{1},p_{1})\in T^{*}Q_{1}$, we have:
    $\pi_{2}\circ\hat{f}^{\lambda}(z_{1})=\pi_{2}(z_{2})=q_{2}=f(q_{1})=f\circ\pi_{1}(z_{1}),$
    so $\pi_{2}\circ\hat{f}^{\lambda}=f\circ\pi_{1}$.
  • By the definition of the pullback, we know:
    $((\hat{f}^{\lambda})^{*}\theta_{Q_{2}})_{z_{1}}=(d\hat{f}^{\lambda})^{\ast}_{z_{1}}(\theta_{Q_{2}})_{(z_{2}},$
    where $(d\hat{f}^{\lambda})^{\ast}_{z_{1}}:T^{*}_{z_{2}}Q_{2}\to T^{*}_{z_{1}}Q_{1}$ is the transpose of the differential $d\hat{f}^{\lambda}_{z_{1}}$.
Now, we can proceed with the calculation:
$((\hat{f}^{\lambda})^{*}\theta_{Q_{2}})_{z_{1}}$ $=(d\hat{f}^{\lambda})^{\ast}_{z_{1}}(\theta_{Q_{2}})_{z_{2}}$
$=(d\hat{f}^{\lambda})^{\ast}_{z_{1}}(d\pi_{2})^{\ast}_{z_{2}}p_{2}\quad\text{(since $(\theta_{Q_{2}})_{z_{2}}=(d\pi_{2})^{\ast}_{z_{2}}p_{2}$)}$
[…]
$=(d(f\circ\pi_{1}))^{\ast}_{z_{1}}p_{2}\quad\text{(since $\pi_{2}\circ\hat{f}^{\lambda}=f\circ\pi_{1}$)}$
[…]
$=(d\pi_{1})^{\ast}_{z_{1}}(\lambda p_{1})\quad\text{(since $(df)^{\ast}_{q_{1}}p_{2}=\lambda p_{1}$)}$
[…]
$=\lambda(\theta_{Q_{1}})_{z_{1}}\quad\text{(since $(\theta_{Q_{1}})_{z_{1}}=(d\pi_{1})^{\ast}_{z_{1}}p_{1}$)}.$
Thus, we conclude that:
$(\hat{f}^{\lambda})^{*}\theta_{Q_{2}}=\lambda\theta_{Q_{1}}.$
@@
Since the exterior derivative commutes with the pull-back, we get
@@ 4 Conformally Symplectic Lie Group Actions
is a flow on $M$, where $\exp:\mathfrak{g}\to G$ is the exponential map. [^1] It is standard to use $\exp(x)$ to denote the exponential map and $e^{x}$ to represent the usual exponential function on real numbers. Since much of this paper concerns the multiplicative group $\mathbb{R}^{+}$, we will alternate between $\exp(x)$ and $e^{x}$ to refer to the exponential map of the group $\mathbb{R}^{+}$, depending on which notation is more convenient in a given context. To each element $\xi\in\mathfrak{g}$, there is an associated vector field on $M$, which we denote by $\xi_{M}$, and call the infinitesimal generator of the action corresponding to $\xi$. This vector field is defined as follows:
@@
We now transition to the specific case where $G=\mathbb{R}^{+}$, the multiplicative group of positive real numbers. This shift is important because much of this paper focuses on the action of $\mathbb{R}^{+}$. In this setting, the Lie algebra $\mathfrak{g}$ is identified with $\mathbb{R}$, and the exponential map corresponds to the usual exponential function on real numbers.
We now introduce some important concepts for this work, namely, conformally symplectic actions and scaling symmetries. The following definition generalizes the notion of a symplectic action.
@@
Let $(M,\omega,\theta)$ be an exact symplectic manifold. An action $\Phi:\mathbb{R}^{+}\times M\to M$ of the multiplicative group $\mathbb{R}^{+}$ on $M$ is said to be conformally symplectic with parameter $c\in\mathbb{R}$ if:
$\Phi_{g}^{*}\omega=g^{c}\omega\quad\text{for all}\quad g\in\mathbb{R}^{+}.$
In the special case when $c=0$, the action is symplectic.
Now, consider the Lie algebra element $\xi\in\mathbb{R}$. The map $\Phi_{e^{\tau\xi}}:M\to M$ defines a flow on $M$. The infinitesimal generator of the action corresponding to $\xi$ is the vector field:
@@
Given the conformally symplectic property $\Phi_{g}^{*}\omega=g^{c}\omega$, we can compute the Lie derivative of $\omega$ with respect to $\xi_{M}$ as follows:
$\mathcal{L}_{\xi_{M}}\omega=\left.\frac{d}{d\tau}\right|_{\tau=0}\Phi_{e^{\tau\xi}}^{*}\omega=\left.\frac{d}{d\tau}\right|_{\tau=0}e^{\tau c\xi}\omega=(c\xi)\omega.$
An action with this property is called infinitesimally conformally symplectic.
The following definition generalizes the concept of an invariant function.
@@
Let $(M,\omega,\theta)$ be an exact symplectic manifold, and let $F:M\to\mathbb{R}$ be a function on $M$. Suppose the action $\Phi:\mathbb{R}^{+}\times M\to M$ of the multiplicative group $\mathbb{R}^{+}$ on $M$ is conformally symplectic with parameter $c$. We say that $F$ is conformally invariant if there exists $b\in\mathbb{R}$ such that:
$\Phi_{g}^{*}F=g^{b}F\quad\text{for all}\quad g\in\mathbb{R}^{+}.$
If $b=0$, the function is simply called invariant.
If $\xi_{M}$ is the infinitesimal generator associated with an element $\xi\in\mathbb{R}$ from the Lie algebra, then the Lie derivative of $F$ with respect to $\xi_{M}$ is:
$\mathcal{L}_{\xi_{M}}F=\left.\frac{d}{d\tau}\right|_{\tau=0}\Phi_{e^{\tau\xi}}^{*}F=\left.\frac{d}{d\tau}\right|_{\tau=0}e^{b\xi\tau}F=(b\xi)F.$
This property is referred to as infinitesimal conformal invariance. It extends the classical notion of infinitesimal invariance (see [16] for the classical definition). In the terminology of [3], a function $F$ with this property is called a scaling function.
Scaling symmetries provide an important extension of Hamiltonian symmetries, where both the symplectic form and the Hamiltonian scale under group actions rather than remaining invariant (as in the case of Hamiltonian symmetries). We formalize this concept in the following definition.
@@
Let $(M,\omega,\theta,H)$ be a Hamiltonian system on an exact symplectic manifold, and let $\Phi:\mathbb{R}^{+}\times M\to M$ be a conformally symplectic action. We say that $\Phi$ is a scaling symmetry if $\Phi$ is conformally symplectic and the Hamiltonian $H$ is conformally invariant. That is, there exist $c,b\in\mathbb{R}$ such that:
$\Phi_{g}^{*}\omega=g^{c}\omega,\quad\Phi_{g}^{*}H=g^{b}H\quad\text{for all}\quad g\in\mathbb{R}^{+}.$
If $\xi_{M}$ is the infinitesimal generator of the scaling symmetry corresponding to the Lie algebra element $\xi\in\mathfrak{g}=\mathbb{R}$, then:
@@
The following proposition establishes a relationship between scaling symmetries and Hamiltonian vector fields. This result, which will be instrumental in our discussion of relative equilibria in Section 6, is analogous to Corollary 2.10 but applies specifically to scaling symmetries rather than diffeomorphisms. Its proof follows a similar approach to that of Proposition 2.9 (see also [22]).
@@
Let $(M,\omega,\theta,H)$ be a Hamiltonian system on an exact symplectic manifold, and let $\Phi:\mathbb{R}^{+}\times M\to M$ be a scaling symmetry. Then:
$\Phi_{g}^{*}X_{H}=g^{b-c}X_{H}.$